# Degenerate scalar scenario of two Higgs doublet model with a complex singlet scalar

### v1

@@ Abstract
We study the two Higgs doublet model with a complex singlet scalar whose imaginary part acts as a dark matter (DM). The scattering of DM and quarks, mediated by three CP-even scalars in this model, is suppressed when masses of CP-even scalars are degenerate; that is called the “degenerate scalar scenario”. Based on this scenario, we show that the strong first-order electroweak phase transition (EWPT) can be achieved without conflicting with constraints from the DM relic density and the direct detection experiments. We also discuss a shift of scalar trilinear coupling from the Standard Model prediction, which could be a test of this model in collider experiments.
@@ I Introduction
Dark matter (DM), whose existence has been suggested by cosmological observations, predicts physics beyond the Standard Model (SM). Among the various DM candidates, weakly interacting massive particles (WIMP) are attractive because their abundance can be explained thermally. (the so-called “WIMP miracle”). However, despite the vigorous search, experiments such as accelerator experiments and DM direct detection experiments have not found the DM signal. In particular, the LZ experiment Aalbers et al. (2022) places very strong constraints on the scattering cross sections of WIMP-DM and nucleons, posing a major challenge for models that include WIMP-DM [^1] The updated results of the LZ experiment, which were presented at the international conference “TeV Particle Astrophysics” held at the University of Chicago in August 2024, impose even stricter constraints..
@@
Possible scenarios consistent with the current direct detection experiments include (1) the mass and interaction of WIMP-DM are decoupled from the SM, and (2) DM-nucleon scattering is suppressed by a built-in mechanism in the model. Since no new particles have been found after the Higgs boson discovery, focusing on the second scenario is natural. One model that realizes this scenario is the complex singlet scalar extension of the Standard Model (CxSM) Barger et al. (2009), which includes a pseudo Nambu-Goldstone (pNG) DM. In the CxSM, as the name suggests, the SM is extended by adding a complex scalar single particle field $S$. The imaginary part of $S$ behaves as the pNG DM, whose stability is guaranteed by the CP symmetry of the scalar potential. On the other hand, the real part of $S$ mixes with the SM Higgs boson to form the mass eigenstates $h_{1}$ and $h_{2}$. These two particles mediate DM-quark scattering. In ref. Abe et al. (2021), it is proposed that when the masses of the two Higgs bosons are degenerate, i.e., $m_{h_{1}}\simeq m_{h_{2}}$, this scattering is suppressed. This suppression mechanism is called the degenerate scalar scenario.
@@
On the other hand, baryon asymmetry of the Universe (BAU), along with DM, is one of the unsolved problems. The most testable scenario to realize BAU is electroweak baryogenesis (EWBG) Kuzmin et al. (1985); Rubakov and Shaposhnikov (1996); Funakubo (1996); Riotto (1998); Trodden (1999); Bernreuther (2002); Cline (2006); Morrissey and Ramsey-Musolf (2012); Konstandin (2013); Senaha (2020) associated with the Higgs physics. EWBG requires strong first-order electroweak phase transition (EWPT), which suggests the need for an extension of the SM Kajantie et al. (1996); Rummukainen et al. (1998); Csikor et al. (1999); Aoki et al. (1999). The CxSM is an attractive model for both DM and BAU, but in ref. Cho et al. (2021), the poor compatibility between the degenerate scalar scenario and strong first-order EWPT driven by the tree-level structure of the effective potential has been pointed out. In order to explain them simultaneously, the DM mass needs to be half the Higgs mass, but here the DM relic abundance is very small due to the resonance effect of the DM annihilation process, and is far from the observed value $\Omega_{\mathrm{DM}}h^{2}=0.1200\pm 0.0012$ Aghanim et al. (2020).
@@
In this paper, we consider a model in which a SU(2)L scalar doublet and a SU(2)L complex scalar singlet are added to the SM. Since this model is the two Higgs doublet model (2HDM) with the complex singlet $S$, we refer to this model as 2HDMS. In the 2HDMS, the CP-even scalars of the two Higgs doublet and the real part of $S$ mix to form three Higgs particles. The imaginary part of $S$ acts as the DM. We investigate the feasibility of the degenerate scalar scenario for four types of Yukawa interaction between the SM fermions and the SU(2)L doublet Higgs bosons, which are distinguished by the $Z_{2}$ charge of each fermion. In the 2HDMS, the one-loop effect is most important for strong first-order EWPT, unlike the CxSM, where the tree-level structure is extremely important. As a result, it is shown that the degenerate scalar scenario is sufficiently valid even in the parameter region where strong first-order EWPT is achieved. In order to resolve the conflict between the degenerate scalar scenario and strong first-order EWPT, a two-component DM model that adds a fermion DM to the CxSM is also being studied Cho et al. (2024), but here we explain the current DM abundance using only the pNG DM of the 2HDMS.
@@ II The Model
where only $m_{3}^{2}$ term breaks the $Z_{2}$ symmetry of the doublets softly. The second term in (1) is given by the singlet $S$ and doublets as
@@
At this time, $v=\sqrt{v_{1}^{2}+v_{2}^{2}}$ = 246.22 GeV and $\tan\beta=v_{2}/v_{1}$ is defined. The following unphysical scalar particles are present in this model; the charged scalar $\quantity(\phi_{i}^{+})$, the CP-odd scalar $\quantity(\eta_{i})$, and the CP-even scalars $\quantity(h_{1},h_{2},s)$. On the other hand, a pseudoscalar $\chi$ is the physical DM candidate.
@@
On the other hand, the mass eigenstates of the charged scalar $\phi^{+}$ and the CP-odd scalar $\eta_{i}$ are defined as follows, respectively:
@@
and $G^{\pm}$ and $G^{0}$ are Nambu-Goldstone bosons. The mass terms of the charged Higgs $H^{\pm}$ and the CP-odd scalars $m_{A}$ are
@@
$m_{A}^{2}$ $=\frac{m_{3}^{2}}{\sin\beta\cos\beta}-\lambda_{5}v^{2}.$ (38)
Also, DM mass is given by
@@
We mention the tree-level flavor changing neutral current (FCNC) restricted experimentally. The Lagrangian of the Yukawa interaction is given as follows:
$\mathcal{L}_{\text{Yukawa }}^{2\mathrm{HDMS}}=-\bar{Q}_{L}Y_{u}\tilde{\Phi}_{u}u_{R}-\bar{Q}_{L}Y_{d}\Phi_{d}d_{R}-\bar{L}_{L}Y_{\ell}\Phi_{\ell}\ell_{R}+\mathrm{h.c.},$ (40)
where $\tilde{\Phi}_{u}\equiv i\sigma_{2}\Phi^{*}_{u}$ and $\sigma_{2}$ being the Pauli matrix. ${Q}_{L}$ and ${L}_{L}$ represent left-handed quarks and left-handed leptons, i.e., $Q_{L}=\left(u_{L},d_{L}\right)^{\top},~{}L_{L}=\left(\nu_{L},e_{L}\right)^{\top}$. $u_{R},~{}d_{R}$, and $l_{R}$ are right-handed up-type quarks, right-handed down-type quarks, and right-handed leptons. $Y_{f}(f=u,d,l)$ is the 3$\times$3 Yukawa matrix of the fermions. Here, in order to prevent the tree-level FCNC, we consider the case where one of the doublets couples with each fermion. As shown in Table 1, by assigning the $Z_{2}$ charge for the fermions, we can determine the combination of ${\Phi}_{f}~{}(f=u,d,l)$ and classify them into four types.
@@
Table 1: Assignment of $Z_{2}$ charge to the fermions and combination of the doublets that couples to each fermion.
@@ III Degenerate scalar scenario
$i\mathcal{M}_{3}$ $=-i2C_{\chi\chi H_{3}}C_{qqH_{3}}\frac{1}{t-m_{H_{3}}^{2}}\bar{u}(p_{4})u(p_{2}),$ (45)
where $t$ describes a momentum transfer, $t\equiv\left(p_{1}-p_{3}\right)^{2}$, and $u(p)~{}(\bar{u}(p))$ represents an incoming (outgoing) quark spinor with a momentum $p$. A factor 2 in the right-hand side of (43), (44) and (45) is a symmetry factor for the $\chi\chi H_{i}$ vertex.
Note that the doublets that couple to each fermion differ depending on the types in the 2HDMS. Therefore, we investigate the suppression mechanism for (A) and (B):
@@ III.1 Type-I, Type-X
$C_{uuH_{i}}=\sum_{i=1}^{3}\frac{m_{u}}{v_{2}}O_{2i},~{}C_{ddH_{i}}=\sum_{i=1}^{3}\frac{m_{d}}{v_{2}}O_{2i},$ (47)
with $m_{u,d}$ being masses of up-type and down-type quarks. On the other hand, the scalar trilinear interaction between $\chi$ and $H_{i}$ is described by
@@
Since the momentum transfer $t$ in the direct detection experiments is very small as compared to the mediator masses, i.e., $t\ll m_{H_{i}}^{2}$, the amplitude (42) can be written by
$i\mathcal{M}$ $=2i\bar{u}\left(p_{4}\right)u\left(p_{2}\right)\frac{\left(m_{u}+m_{d}\right)}{v_{2}}\frac{C_{\chi\chi H_{i}}O_{2i}}{m_{H_{i}}^{2}},$
$=2i\bar{u}\left(p_{4}\right)u\left(p_{2}\right)\frac{\left(m_{u}+m_{d}\right)}{v_{2}}\left(C_{\chi\chi h_{1}}\frac{O_{1i}O_{2i}}{m_{H_{i}}^{2}}+C_{\chi\chi{h_{2}}}\frac{O_{2i}^{2}}{m_{H_{i}}^{2}}+C_{\chi\chi s}\frac{O_{3i}O_{2i}}{m_{H_{i}}^{2}}\right).$ (50)
Therefore, due to the orthogonality of the mixing matrix (32), the amplitude vanishes in the degenerate scalar scenario $(m_{H_{1}}=m_{H_{2}}=m_{H_{3}})$.
III.2 Type-II, Type-Y
The Yukawa interaction for Type-II and Type-Y is given by
$-\mathcal{L}_{\mathrm{Yukawa}}$ $=\sum_{i=1}^{3}C_{uuH_{i}}\overline{u_{L}}u_{R}H_{i}+\sum_{i=1}^{3}C_{ddH_{i}}\overline{d_{L}}d_{R}H_{i},$ (51)
@@ III.2 Type-II, Type-Y
$C_{uuH_{i}}=\sum_{i=1}^{3}\frac{m_{u}}{v_{2}}O_{2i},~{}C_{ddH_{i}}=\sum_{i=1}^{3}\frac{m_{d}}{v_{1}}O_{1i}.$ (52)
The amplitude (42) becomes
$i\mathcal{M}$ $=2i\bar{u}\left(p_{4}\right)u\left(p_{2}\right)\left(\frac{m_{u}}{v_{2}}\frac{C_{\chi\chi H_{i}}O_{2i}}{m_{H_{i}}^{2}}+\frac{m_{d}}{v_{1}}\frac{C_{\chi\chi H_{i}}O_{1i}}{m_{H_{i}}^{2}}\right)$
$=2i\bar{u}\left(p_{4}\right)u\left(p_{2}\right)\frac{m_{u}}{v_{2}}\left(C_{\chi\chi h_{1}}\frac{O_{1i}O_{2i}}{m_{H_{i}}^{2}}+C_{\chi\chi{h_{2}}}\frac{O_{2i}^{2}}{m_{H_{i}}^{2}}+C_{\chi\chi s}\frac{O_{3i}O_{2i}}{m_{H_{i}}^{2}}\right)$
$+2i\bar{u}\left(p_{4}\right)u\left(p_{2}\right)\frac{m_{d}}{v_{1}}\left(C_{\chi\chi h_{1}}\frac{O_{1i}^{2}}{m_{H_{i}}^{2}}+C_{\chi\chi{h_{2}}}\frac{O_{2i}O_{1i}}{m_{H_{i}}^{2}}+C_{\chi\chi s}\frac{O_{3i}O_{1i}}{m_{H_{i}}^{2}}\right).$ (53)
As for case A (Type-I, Type-X), this scattering is suppressed when the three Higgs masses are equal.
We comment on the sum rule for the degenerate scalar scenario taking case A (Type-I, Type-X) as an example. The scattering of DM and quarks for case A (Type-I, Type-X) is as shown in Fig. 1. The left side uses the mass eigenstates, while the right side uses the gauge eigenstates.
Figure 1: Feynman diagrams of the DM-quark scattering in the mass and gauge eigenstates of scalar mediators.
In order for the scalar $s$ to couple with the quarks $q$, the scalar $s$ must be converted into $h_{2}$ through the coupling $C_{hs}$ (the right diagram in Fig. 1). On the other hand, the amplitude mediated by $h_{2}$ is proportional to the coupling $C_{\chi\chi h_{2}}$ (the middle diagram in Fig. 1). Therefore, the amplitudes are canceled only when $C_{hs}$ and $C_{\chi\chi h_{2}}$ satisfy a specific relationship.
@@
where parameters $A,\Delta_{1},\Delta_{2}$ and $\Delta_{3}$ are given by parameters besides $C_{h_{1}s},C_{h_{2}s}$ and $C_{ss}$ in the scalar potential. Extracting the part that is involved in suppressing the scattering, one gets
@@
$=\frac{A}{v_{s}m_{H_{i}}^{2}}\left(O_{2i}O_{3i}m_{H_{i}}^{2}+\Delta_{1}O_{1_{i}}O_{2_{i}}+\Delta_{2}O_{2_{i}}^{2}+\Delta_{3}O_{3_{i}}O_{2_{i}}\right).$ (55)
Therefore, the amplitude in the degenerate Higgs region vanishes when
@@
$C_{\chi\chi h_{2}}$ $=\frac{\delta_{2}}{4}v_{2},$ (58)
$C_{h_{2}s}$ $=\frac{\delta_{2}}{2}v_{2}v_{s}.$ (59)
Comparing (58) and (59) with (54), we find

### v2

@@ Abstract
We study the two Higgs doublet model with a complex singlet scalar whose imaginary part acts as dark matter (DM). The scattering of DM and quarks, mediated by three CP-even scalars in this model, is suppressed when masses of CP-even scalars are degenerate; that is called the “degenerate scalar scenario”. Based on this scenario, we show that the strong first-order electroweak phase transition (EWPT) can be achieved without conflicting with constraints from the DM relic density and the direct detection experiments. We also discuss a shift of scalar trilinear coupling from the Standard Model prediction, which could be a test of this model in collider experiments.
@@ I Introduction
Dark matter (DM), whose existence has been suggested by cosmological observations, predicts physics beyond the Standard Model (SM). Among the various DM candidates, weakly interacting massive particles (WIMP) are attractive because their abundance can be explained thermally. It is so-called the“WIMP miracle”. However, despite the vigorous searches, such as accelerator experiments and DM direct detection experiments have not found the DM signal. In particular, the LZ experiment Aalbers et al. (2024) places very strong constraints on the scattering cross sections of WIMP-DM and nucleons, posing a major challenge for models that include WIMP-DM.
@@
Possible scenarios consistent with the current direct detection experiments include (1) DM is either sufficiently massive or has only tiny interactions with the SM particles and (2) DM-nucleon scattering is suppressed by a built-in mechanism in the model. Here, we focus on the second scenario. One model that realizes this scenario is the complex singlet scalar extension of the Standard Model (CxSM) Barger et al. (2009), which includes a pseudo Nambu-Goldstone (pNG) DM. In the CxSM, as the name suggests, the SM is extended by adding a complex scalar single particle field $S$. The imaginary part of $S$ behaves as the pNG DM, whose stability is guaranteed by the CP symmetry of the scalar potential. On the other hand, the real part of $S$ mixes with the SM Higgs boson to form the mass eigenstates $h_{1}$ and $h_{2}$. These two particles mediate DM-quark scattering. In ref. Gross et al. (2017), the DM-quark scattering is suppressed due to the small momentum transfer. However, this model has a so-called domain wall problem because the $Z_{2}$ symmetry ($S\to-S$) of the scalar potential $V$ is spontaneously broken due to the development of the vacuum expectation value (VEV) of the singlet scalar $(S)$. Therefore, in ref. Abe et al. (2021), a linear term of $S$ ($V\supset a_{1}S$) that breaks the $Z_{2}$ symmetry is introduced, and in this case, it is proposed that when the masses of the two Higgs bosons are degenerate, i.e., $m_{h_{1}}\simeq m_{h_{2}}$, this scattering is suppressed. This suppression mechanism is called the degenerate scalar scenario.
@@
On the other hand, baryon asymmetry of the Universe (BAU), along with DM, is one of the unsolved problems. The most testable scenario to realize BAU is electroweak baryogenesis (EWBG) Kuzmin et al. (1985); Rubakov and Shaposhnikov (1996); Funakubo (1996); Riotto (1998); Trodden (1999); Bernreuther (2002); Cline (2006); Morrissey and Ramsey-Musolf (2012); Konstandin (2013); Senaha (2020) associated with the Higgs physics. EWBG requires strong first-order electroweak phase transition (EWPT), which suggests the need for an extension of the SM Kajantie et al. (1996); Rummukainen et al. (1998); Csikor et al. (1999); Aoki et al. (1999). The CxSM is an attractive model for both DM and BAU, but in ref. Cho et al. (2021), the poor compatibility between the degenerate scalar scenario and strong first-order EWPT. In almost all models, strong first-order EWPT is caused by the effects of the finite temperature and the quantum corrections on the scalar potential, the latter is known as the Coleman-Weinberg potential. However, in some models with extended Higgs sector, the potential barrier is enhanced by mixing between the electroweak Higgs sector and a new scalar sector at the tree level. As investigated in ref. Cho et al. (2021), the contribution of the latter is large in the CxSM, and a large $\delta_{2}$, which is the mixing parameter between the SM Higgs and the singlet scalar $S$, is favored to realize strong first-order EWPT. On the other hand, a small $\delta_{2}$ achieved through the degenerate scalar scenario suppresses the DM-quark scattering, so these two give contradictory conditions for the parameters. To explain them simultaneously, the DM mass needs to be half the Higgs mass, but here, the DM relic abundance is very small due to the resonance effect of the DM annihilation process and is far from the observed value $\Omega_{\mathrm{DM}}h^{2}=0.1200\pm 0.0012$ Aghanim et al. (2020).
@@
In this paper, we consider a model in which an SU(2)L doublet scalar and a complex singlet scalar are added to the SM. Since this model is the two-Higgs doublet model (2HDM) with the complex singlet $S$, we refer to this model as 2HDMS. In the 2HDMS, the CP-even scalars of the two Higgs doublet and the real part of $S$ mix to form three Higgs particles $H_{1,2,3}$. The imaginary part of $S$ acts as the DM. There are studies on 2HDMS, such as Jiang et al. (2019); Zhang et al. (2021); Biekötter and Olea-Romacho (2021); Biekötter et al. (2022), but in these models, $S$ has the $Z_{2}$ symmetry and may suffer from the domain wall problem, so here we introduce a linear term for $S$ that breaks the symmetry. Furthermore, since $v_{S}$ becomes nonzero for any temperature due to $a_{1}\neq 0$, the EWPT is also different from the model with $a_{1}=0$. As with the CxSM mentioned earlier, we investigate whether the DM-quark scattering is suppressed when the mediator’s masses are degenerate ($m_{H_{1}}\simeq m_{H_{2}}\simeq m_{H_{3}}$), in other words, we investigate the feasibility of the degenerate scalar scenario. We take into account four types of Yukawa interaction between the SM fermions and the SU(2)L doublet Higgs bosons, which are distinguished by the $Z_{2}$ charge of each fermion. In the 2HDMS, the one-loop effect is most important for strong first-order EWPT, unlike the CxSM, where the tree-level structure is extremely important. As a result, it is shown that the degenerate scalar scenario is sufficiently valid even in the parameter region where strong first-order EWPT is achieved. In order to resolve the conflict between the degenerate scalar scenario and strong first-order EWPT, a two-component DM model that adds a fermion DM to the CxSM is also being studied Cho et al. (2024), but here we explain the current DM abundance using only the pNG DM of the 2HDMS.
@@ II The Model
where only $m_{3}^{2}$ term breaks the $Z_{2}$ symmetry of the doublets ($\Phi_{1}\to+\Phi_{1},~{}\Phi_{2}\to-\Phi_{2}$) softly. The second term in (1) is given by the singlet $S$ and doublets as
@@
At this time, $v=\sqrt{v_{1}^{2}+v_{2}^{2}}$ = 246.22 GeV and $\tan\beta=v_{2}/v_{1}$ is defined. The following scalar particles are present in this model; the charged scalar $\quantity(\phi_{i}^{+})$, the CP-odd scalar $\quantity(\eta_{i})$, and the CP-even scalars $\quantity(h_{1},h_{2},s)$. On the other hand, a pseudoscalar $\chi$ is the physical DM candidate.
@@
On the other hand, the charged Higgs $H^{\pm}$ and the CP-odd scalar $A$ are defined from $\phi^{+}_{1,2},~{}\eta_{1,2}$ as
@@
In (33), $G^{\pm}$ and $G^{0}$ denote Nambu-Goldstone bosons. The mass eigenvalues of $H^{\pm}$ and $A$ are given by
@@
$m_{A}^{2}$ $=\frac{m_{3}^{2}}{\sin\beta\cos\beta}-\lambda_{5}v^{2},$ (38)
respectively. Also, DM mass is given by
@@
In order to prevent the tree-level FCNC in this model, we consider the case where one of the doublets couples with each fermion Branco et al. (2012). The Lagrangian of the Yukawa interaction is given as follows:
$-\mathcal{L}_{\text{Yukawa }}=\bar{Q}_{L}Y_{u}\tilde{\Phi}_{u}u_{R}+\bar{Q}_{L}Y_{d}\Phi_{d}d_{R}+\bar{L}_{L}Y_{\ell}\Phi_{\ell}\ell_{R}+\mathrm{h.c.},$ (40)
where $\Phi_{f}(f=u,d,\ell)$ is a Higgs doublet that couples to the fermion $f$ and $\tilde{\Phi}_{u}\equiv i\sigma_{2}\Phi^{*}_{u}$ and $\sigma_{2}$ being the Pauli matrix. ${Q}_{L}$ and ${L}_{L}$ represent left-handed quarks and left-handed leptons, i.e., $Q_{L}=\left(u_{L},d_{L}\right)^{\top},~{}L_{L}=\left(\nu_{L},e_{L}\right)^{\top}$. $u_{R},~{}d_{R}$, and $\ell_{R}$ are right-handed up-type quarks, right-handed down-type quarks, and right-handed leptons. $Y_{f}(f=u,d,\ell)$ is the 3$\times$3 Yukawa matrix of the fermions. As shown in Table 1, by assigning the $Z_{2}$ charge for the fermions, we can determine the combination of ${\Phi}_{f}~{}(f=u,d,\ell)$ and classify them into four types.
@@
Table 1: Assignment of $Z_{2}$ charge to the fermions and combination of the doublets that couples to each fermion. The Higgs doublets $\Phi_{1}$ and $\Phi_{2}$ transform as $\Phi_{1}\to+\Phi_{1}$ and $\Phi_{2}\to-\Phi_{2}$, respectively.
@@ III Degenerate scalar scenario
$i\mathcal{M}_{3}$ $=-i2C_{\chi\chi H_{3}}C_{qqH_{3}}\frac{1}{t-m_{H_{3}}^{2}}\bar{u}(p_{4})u(p_{2}).$ (45)
Here, $C_{S_{1}S_{2}S_{3}}$ denotes the scalar trilinear coupling for $S_{1},S_{2}$ and $S_{3}$, i.e., $\mathcal{L}\supset C_{S_{1}S_{2}S_{3}}S_{1}S_{2}S_{3}$. Thus, $C_{\chi\chi H_{i}}$ is the three-point scalar coupling between two $\chi$ and $H_{i}$ ($i=1,2,3$). On the other hand, $C_{qqH_{i}}$ is the Yukawa coupling between the SM quarks $q$ and $H_{i}$, $\mathcal{L}\supset C_{qqH_{i}}\bar{q}_{L}q_{R}H_{i}$. A variable $t$ describes a momentum transfer, $t\equiv\left(p_{1}-p_{3}\right)^{2}$, and $u(p)~{}(\bar{u}(p))$ represents an incoming (outgoing) quark spinor with a momentum $p$. A factor 2 in the right-hand side of (43), (44) and (45) is a symmetry factor for the $\chi\chi H_{i}$ vertex.
First, we evaluate the degenerate scalar scenario in the mass eigenstates, and then use the gauge eigenstate to understand this scenario better. In the 2HDMS, the scalar trilinear couplings are given as follows:
$C_{\chi\chi H_{i}}=\frac{1}{2v_{S}}\left(m_{H_{i}}^{2}+\frac{\sqrt{2}a_{1}}{v_{S}}\right)O_{3i},$ (46)
which is derived from Eq. (17). On the other hand, the Yukawa interactions can be expressed as
$C_{qqH_{i}}=\left\{\begin{array}[]{l}\frac{m_{q}}{v_{1}}O_{1i}\quad\text{(When $q$ couples to $\Phi_{1}$)},\\ \frac{m_{q}}{v_{2}}O_{2i}\quad\text{(When $q$ couples to $\Phi_{2}$)},\end{array}\right.$ (49)
where $m_{q}$ is the mass of quark $q$. Therefore, the amplitude (42) becomes
$i\mathcal{M}=\left\{\begin{array}[]{l}-i2\frac{1}{2v_{S}}\sum_{i=1}^{3}\frac{m_{H_{i}}^{2}+\frac{\sqrt{2}a_{1}}{v_{S}}}{t-m_{H_{i}}^{2}}O_{3i}O_{1i}\bar{u}(p_{4})u(p_{2})\quad\text{(When $q$ couples to $\Phi_{1}$)},\\ -i2\frac{1}{2v_{S}}\sum_{i=1}^{3}\frac{m_{H_{i}}^{2}+\frac{\sqrt{2}a_{1}}{v_{S}}}{t-m_{H_{i}}^{2}}O_{3i}O_{2i}\bar{u}(p_{4})u(p_{2})\quad\text{(When $q$ couples to $\Phi_{2}$)}.\end{array}\right.$ (52)
If $a_{1}=0$, the amplitude vanishes in the low-energy limit ($t\to 0$) due to the orthogonality of the mixing matrix (32). However, as mentioned above, in order to avoid the domain wall problem associated with the development of $v_{S}$, we are currently setting $a_{1}$ to a nonzero value. In this case, the Higgs masses $m_{H_{i}}$ must be degenerate to suppress the amplitude. This is the mechanism we call the degenerate scalar scenario.
Next, we discuss the degenerate scalar scenario in the gauge eigenstates to explore the origin of this mechanism. The doublets that couple to each fermion differ depending on the types in the 2HDMS. Therefore, we investigate the suppression mechanism for (A) and (B):
@@ III.1 Type-I, Type-X
$C_{uuH_{i}}=\frac{m_{u}}{v_{2}}O_{2i},~{}C_{ddH_{i}}=\frac{m_{d}}{v_{2}}O_{2i},$ (54)
with $m_{u,d}$ being masses of up-type and down-type quarks. Here, the lepton sector is omitted since we focus only on the scattering with quarks. On the other hand, the scalar trilinear interaction between $\chi$ and $H_{i}$ is described by
@@
Explicit expression of these couplings are summarized in Appendix B. Since the momentum transfer $t$ in the direct detection experiments is very small as compared to the mediator masses, i.e., $t\ll m_{H_{i}}^{2}$, the amplitudes related to up-type quarks and down-type quarks can each be written by
$i\mathcal{M}_{\mathrm{up}}$ $=2i\bar{u}\left(p_{4}\right)u\left(p_{2}\right)\frac{m_{u}}{v_{2}}\sum_{i=1}^{3}\frac{C_{\chi\chi H_{i}}O_{2i}}{m_{H_{i}}^{2}},$
$=2i\bar{u}\left(p_{4}\right)u\left(p_{2}\right)\frac{m_{u}}{v_{2}}\sum_{i=1}^{3}\left(C_{\chi\chi h_{1}}\frac{O_{1i}O_{2i}}{m_{H_{i}}^{2}}+C_{\chi\chi{h_{2}}}\frac{O_{2i}^{2}}{m_{H_{i}}^{2}}+C_{\chi\chi s}\frac{O_{3i}O_{2i}}{m_{H_{i}}^{2}}\right),$ (57)
$i\mathcal{M}_{\mathrm{down}}$ $=2i\bar{u}\left(p_{4}\right)u\left(p_{2}\right)\frac{m_{d}}{v_{2}}\sum_{i=1}^{3}\frac{C_{\chi\chi H_{i}}O_{2i}}{m_{H_{i}}^{2}},$
$=2i\bar{u}\left(p_{4}\right)u\left(p_{2}\right)\frac{m_{d}}{v_{2}}\sum_{i=1}^{3}\left(C_{\chi\chi h_{1}}\frac{O_{1i}O_{2i}}{m_{H_{i}}^{2}}+C_{\chi\chi{h_{2}}}\frac{O_{2i}^{2}}{m_{H_{i}}^{2}}+C_{\chi\chi s}\frac{O_{3i}O_{2i}}{m_{H_{i}}^{2}}\right).$ (58)
Therefore, when three Higgs masses are degenerate, due to the orthogonality of the mixing matrix (32), the terms that are proportional to $O_{1i}O_{2i}$ and $O_{3i}O_{2i}$ vanish. On the other hand, the term that is proportional to $O_{2i}^{2}$ includes $C_{\chi\chi{h_{2}}}$. In this model, $C_{\chi\chi{h_{2}}}$ can be expressed as
$C_{\chi\chi{h_{2}}}=\frac{\delta_{2}}{4}v_{2},$ (59)
@@
$\delta_{2}$ $=\frac{2}{v_{2}v_{S}}\sum_{i=1}^{3}O_{2i}O_{3i}m_{H_{i}}^{2}.$ (60)
$\delta_{2}$ is also suppressed by the orthogonality of the mixing matrix. Therefore, the amplitudes vanish in the degenerate scalar scenario $(m_{H_{1}}=m_{H_{2}}=m_{H_{3}})$. However, note that this suppression mechanism may not work well when $v_{2}$ and $v_{S}$, which appear in the denominator of Eq. (60), are very small.
We comment on the sum rule for the degenerate scalar scenario. The scattering of DM and quarks for case A (Type-I, Type-X) is as shown in Fig. 1. The left side uses the mass eigenstates, while the right side uses the gauge eigenstates.
Figure 1: Feynman diagrams of the DM-quark scattering in the mass and gauge eigenstates of scalar mediators. Both up-type quarks and down-type quarks couple to $\Phi_{2}$ in case A (Type-I, Type-X).
In order for the scalar $h_{1}$ and $s$ to couple with the quarks $q$, $h_{1}$ and $s$ must be converted into $h_{2}$ through the couplings $C_{h_{1}h_{2}}$ and $C_{h_{2}s}$ (the left and right diagram of the right-handed side in Fig. 1). On the other hand, the amplitude mediated by $h_{2}$ is proportional to the coupling $C_{\chi\chi h_{2}}$ (the middle of the right-handed side diagram in Fig. 1).
@@
where parameters $A,\Delta_{1},\Delta_{2}$ and $\Delta_{3}$ are expressed by parameters besides $C_{h_{1}s},C_{h_{2}s}$ and $C_{ss}$ in the scalar potential. The bilinear couplings are also given in Appendix B. Referring to Eqs. (57), (58) and extracting the part that is involved in suppressing the scattering, one gets
@@
$=\frac{A}{v_{S}m_{H_{i}}^{2}}\left(O_{3i}O_{2i}m_{H_{i}}^{2}+\Delta_{1}O_{1i}O_{2i}+\Delta_{2}O_{2i}^{2}+\Delta_{3}O_{3i}O_{2i}\right),$ (62)
where we use relations of coefficients between the gauge and mass eigenstates, i.e., $C_{h_{1}s}O_{1i}+C_{h_{2}s}O_{2i}+C_{ss}O_{3i}=O_{3i}m_{H_{i}}^{2}$. [^1] This formula is derived from $O^{\top}M_{S}^{2}O=\sum_{i=1}^{3}m_{H_{i}}^{2}$ related to Eq. (17). Therefore, the amplitude in the degenerate Higgs region vanishes when
@@
$C_{\chi\chi h_{2}}=\frac{\delta_{2}}{4}v_{2},\quad C_{h_{2}s}=\frac{\delta_{2}}{2}v_{2}v_{S}.$ (65)
Comparing (65) with (61), we find